Given TanA=1/2,tanB=1/5,tanC=1/8 show that A+B+C=π/2

 Given
TanA=1/2,tanB=1/5,tanC=1/8 show that 
A+B+C=π/2
 

Grade:12th pass

1 Answers

kkbisht
5 years ago
Please check your Question: The answer is :A+B+C=pie/4
use the Trigonometric formula : Tan(A+B+C) = (Tan A + TanB + TanC- Tan A TanB TanC)/(1-TanATanB- TanBTanC-TanCTanA)
Now substitute the values given for : TanA=1/2 , TanB=1/5, TanC=1/8
Weg get:Tan(A+B+C)= (1/2 +1/5+1/8-  1/2 .1/5. 1/8)/(1-1/2.1/5 -1/5.1/8-1/8.1/5) 
                                   = 65/65= 1 => A+B+C= (pie)/4
kkbisht